OpenStudy (anonymous):

Choose the equation of the line passing through the point (3, -2) and parallel to y = 3x - 6.

OpenStudy (anonymous):

y+2 = 3(x-3) point slope form

OpenStudy (anonymous):

okay so now its in point slope form. but i still can't comprehend how i get the new y value from this step.

OpenStudy (anonymous):

i always get something that's not the answer.

OpenStudy (anonymous):

-.- try simplifying that

OpenStudy (anonymous):

y+2=3x-9

OpenStudy (anonymous):

then..

OpenStudy (anonymous):

wat's the answer?

OpenStudy (anonymous):

oh okay im over thinking it mybad.

OpenStudy (anonymous):

thanks